OpenStudy (anonymous):

Select the factors of xy - 4x + 2y - 8.

OpenStudy (anonymous):

optins..??

OpenStudy (aum):

xy - 4x + 2y - 8 = x(y - 4) + 2(y - 4) = (x + 2)(y - 4)